OpenStudy (anonymous):

Write y = x + 5 in standard form using integers.

OpenStudy (aravindg):

take all terms to one side

OpenStudy (anonymous):

A. –x – 6y = 30 B. 6x – y = 30 C. –x + 6y = 30 D. –x + 6y = 5